OpenStudy (anonymous):

what is the raduis of cylindrical object if it height is 10 m and a volume of 1538.6 cubic meter?

OpenStudy (anonymous):

\[r = \sqrt(V/\pi*H)\] just put H = 10, and V that horrible number..and take pi ~ 3.14

OpenStudy (anonymous):

because \[V = BH, B = r^2H \rightarrow V = r^2 \pi H\]

OpenStudy (anonymous):

B = r^2 * pi, sorry

OpenStudy (anonymous):

is there a simple way? im only 2nd year

OpenStudy (anonymous):

ok, I'll do the whole procedure :) just a sec :)

OpenStudy (anonymous):

Fist, basic expression for cylinder's volume is: V = B * H V - volume, given, 1538.6 B - are of circle in the base H - height, here 10m Since base is a circle, formula for are of circle is: \[B = r^2 \pi\] just put that together.. \[V = B H = r^2 \pi H\] now, lets replace numbers \[1538.6 = r^2 \pi * 10\] put 10 and \[\pi\] on the other side of "=" (if its up on one side, it goes down on the other) \[r^2 = V / (\pi H)\] \[r^2 = 1538.6 / (3.14 * 10)\] \[r^2 = 1538.6 / 31.4\] so, calculate that number...and r is square root of that number :) there just isn't any simpler way :)

OpenStudy (anonymous):

am, ok... and there is 7 remaining questions, cn u help me?